# Float Self-Tagging

### v1

@@ Abstract.
Dynamic and polymorphic languages must attach information, such as types, to run time objects, and therefore adapt the memory layout of values to include space for this information. This is especially problematic in the case of IEEE754 double-precision floating-point numbers, which require exactly 64 bits, leaving no space for type information. The two main encodings in-use to this day, tagged pointers and NaN-tagging, either allocate floats on the heap or unbox them at the cost of an overhead when handling all other objects.
This paper presents self-tagging, a new approach to object tagging that can attach type information to 64-bit objects while retaining the ability to use all of their 64 bits for data. At its core, self-tagging exploits the fact that some bit sequences appear with very high probability. Superimposing tags with these frequent sequences allows encoding both 64-bit data and type within a single machine word. Implementations of self-tagging demonstrate that it unboxes all floats in practice, accelerating the execution time of float-intensive benchmarks in Scheme by 2.3$\times$, and in JavaScript by 2.7$\times$ without impacting the performance of other benchmarks, which makes it a good alternative to both tagged pointers and NaN-tagging.
Dynamic Languages, Polymorphic Languages, Floating-Point, NaN-Tagging, Compiler Optimization, JavaScript, Scheme
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Source code generationccs: Software and its engineering Polymorphism
@@ 1. Introduction
In dynamic and other polymorphic languages, efficiently handling floating-point numbers (floats) remains an elusive problem in search of an optimal solution. These languages require attaching type information to values at run time, including numeric types like integers and floats. This requirement conflicts with the IEEE754 standard (iee, 2019) for encoding floats, which uses a full 64-bit word to represent a double-precision float, leaving no space for additional type information. Various representations have been developed to mitigate this issue, each with its own trade-offs.
This paper presents a novel technique for encoding floats that avoids most float allocations without adversely affecting the cost of checking types and accessing objects, which are frequent operations of dynamic languages. This technique does not require any programs static analysis and could be integrated in most runtime systems of languages that need run time types. It is developed in the context of 64-bit runtime systems but also has applications on 32-bit systems. The rest of this section presents the main encodings in-use as well as their trade-offs.
@@ 1.1. Tagged Objects
The classical and popular solution to preserve type information is to attach a tag to all objects (Gudeman, 1993; Leijen, 2022; Jones et al., 2012a; Appel, 1998). A machine word is used to encode an object reference (either a value or a pointer to a heap-allocated value) and a short sequence of bits is reserved in the word to store type information. For the rest of the paper the term object is used interchangeably with object reference. Tagging allows low-cost type checks and object access at the cost of losing a few bits for data. Aligning all heap-allocated values to 64-bit machine words conveniently frees the low bits of pointers to store a 3-bit tag. On many architectures, accessing the fields of the object can be done at no additional cost by using an offset in the memory dereference instruction to take into account the tag for that type of object. Objects can thus be encoded using one of the three following representations.
@@
  • Generic pointers represent all remaining object types as pointers with a generic tag. Since the generic tag is associated with multiple types, type information is stored in memory, typically in a header.
Hence, a tag indicates a type, but also a memory representation. A tagged value representation is the most efficient as it needs no heap allocation and no memory read. A tagged pointer still allows efficient type checks, but requires heap allocation. This introduces the overhead of dereferencing the pointer, and adds a strain on the garbage collector. A generic pointer is the least efficient representation as it involves storing type information in memory, which increases space usage and the cost of type checks and memory management. Consequently, specific tags are conscientiously assigned to the most frequent types observed in programs, such as small integers or floats. Figure 1 shows a typical memory layout of each representation.
@@
As an example, consider an implementation using 3-bit tags. Due to the frequent use of small integers, a tagged value representation is appropriate for small integers and the choice of the tag 000 allows addition/subtraction directly on the tagged values. Since 3 bits are occupied by the tag, small integers are limited to 61 bits. This is generally considered an acceptable trade-off since a two’s complement representation of signed integers can accommodate any number of bits, and 61 bits still offers a wide range of values. When larger integers are required, generic pointers to heap-allocated big integers can be used (Leijen, 2022). Consequently this provides a hybrid representation where the most common integers appear as efficient tagged values and less-frequent big integers are represented with more costly generic pointers.
Unfortunately, when it comes to the representation of floats, none of the aforementioned options is well-suited. Double-precision floats cannot be represented as tagged values since the IEEE754 standard enforces the use of 1 sign bit, 11 exponent bits and 52 mantissa bits (see Figure 2(a)), totaling 64 bits and leaving no space for a tag (iee, 2019). Parting from this standard is impracticable due to the set of instructions offered by X86 (Corporation, 2024) and ARM (Limited, 2009) architectures being specific to 32 or 64-bit floats.
@@
Still, tagged objects are straightforward to implement and are thus found in numerous compilers such as V8 (Sheludko and Solanes, 2020), QuickJS (Bellard and Gordon, 2024), and Hopc (Serrano, 2021) (JavaScript), the Lua interpreter (Roberto Ierusalimschy, 2024) (Lua), CRuby (Shaughnessy, 2013) (Ruby), SBCL (Newman et al., 2024) (Lisp), and Bigloo (Manuel Serrano, 2024) (Scheme). Recently, CPython also moved toward tagged pointers in the process of removing its global interpreter lock (Gross, 2023).
1.2. NaN-tagging
NaN-tagging circumvents the drawbacks of tagged pointer floats by reclaiming unused bits in the encodings of floating-point NaN values to store data (Gudeman, 1993; Leijen, 2022). As per the IEEE754 standard, a NaN value is represented by setting all 11 bits of the exponent to 1 and a non-zero mantissa (zero is used to encode Infinity). The mantissa’s highest bit distinguishes between a quiet and signaling NaN, which determines whether the NaN should signal an exception or fall through operations.
A subset of NaN’s can thus be reserved to encode non-float objects. A convenient subset is that of negative, quiet NaN’s, illustrated in Figure 2(b), which correspond to the interval from 0xfff8:$0$:$0$:$0$ to 0xffff:ffff:ffff:ffff. This partitions floating-point numbers in two intervals. Sequences of bits above 0xfff8:$0$:$0$:$0$ are reserved for non-float objects and those lesser than or equal to 0xfff8:$0$:$0$:$0$ are reserved for floats with 0xfff8:$0$:$0$:$0$ itself representing a negative, quiet NaN.
@@ 1.2. NaN-tagging
(b) NaN-tagging uses the 51 free bits on negative, quiet NaN values to store data.
@@
On current hardware, memory addresses typically fit in 48 bits. The 51 bits uncovered by NaN-tagging are thus sufficient for storing tagged pointers with 3-bit tags, and 32-bit small integers. It offers the advantage of unboxed floats, but impacts the performance of all other, non-float objects due to higher-cost machine instructions necessary to check the type and dereference NaN-tagged pointers.
This reliance on hardware specific details also interferes with other optimizations and portability. For instance, NaN-tagging prevents (or largely complicates) optimizations relying on stack allocations. The stack uses high memory addresses that do not fit in 48 bits unless encoded relative to the location of the stack segment. Another caveat of NaN-tagging is the impossibility to apply it on 32-bit architectures as 32-bit NaN’s would severely cripple memory management by leaving only 22-bit pointers, which only allows addressing about 4 MB.
NaN-tagging is used in a few language implementations, including SpiderMonkey (Foundation, 2024) (JavaScript), tinylisp (Engelen, 2024) (Lisp), LuaJIT (Pall, 2024) (Lua), and Zag (Mason, 2022) (Smalltalk).
1.3. Pointer-biased NaN-tagging
NuN-tagging is a variant of NaN-tagging that alleviates the cost of dereferencing NaN-tagged pointers (Leijen, 2022). It relies on the (currently valid) fact that common hardware never returns negative, quiet NaN’s greater than fffd:ffff:ffff:ffff.
Figure 3. Encoding of Nun-tagged values, after applying the 0x0001:$0$:$0$:$0$ offset. The low bit sequences are used for non-float objects that can be represented as tagged pointers or value.
Instead of taking advantage of the unused range of bit sequences from 0xfffe:$0$:$0$:$0$ to 0xffff:ffff:ffff:ffff, which would result in an encoding similar to standard NaN-tagging, NuN-tagging offsets all floats by adding 0x0001:$0$:$0$:$0$. This frees up the lower bit sequences for 48-bit tagged objects as shown in Figure 3. It however reintroduces a cost for encoding and decoding floats to operate on their value.
NuN-tagging is used by JavaScriptCore (Inc., 2024) (JavaScript).
@@ 1.4. Contribution
The contribution of this paper is a new approach to object tagging, named self-tagging, that allows attaching type information to some 64-bit objects while retaining the ability to use all of their 64 bits for storing data. This approach is applied to implement double-precision floats as tagged values instead of tagged pointers. Contrary to NaN-tagging, the newly obtained representation does not impact the performance of encoding and decoding other tagged objects and it is compatible with stack addresses. It is also applicable in the context of single-precision floats on 32-bit architectures, which NaN-tagging does not support. In practice, self-tagging is shown to prevent the heap-allocation of all floats, thus providing the advantages of both tagging and NaN-tagging without their main drawbacks.
@@ 1.5. Paper Structure
This paper is structured as follows. Section 2 describes the general idea of self-tagging with a focus on its application to floats. Section 3 provides a first implementation of self-tagging in the Bigloo Scheme compiler. Section 4 shows how to refine the technique to unbox all floats in the Bigloo compiler and in the Hopc ahead-of-time JavaScript compiler. Section 5 presents technical details of self-tagging implementation. It shows portable implementations that could be used by compilers that generate either C or assembly code. Section 6 discusses the adaptation of self-tagging to 32-bit architectures. Finally, section 7 presents related work.
@@ 2. Self-Tagging
This section describes self-tagging, a new tagging technique that exploits the fact that some values naturally contain the appropriate tag corresponding to their type, at the correct location, in their bit arrangement. While this may appear coincidental, it occurs frequently enough to allow the transformation of a significant proportion of heap allocated objects (either represented as tagged pointers or generic pointers) to tagged values. In this paper, unless stated otherwise, double-precision floats and a 64-bit architecture are implied. However, self-tagging pertains to 32-bit architectures as well, which is discussed in Section 6.
Consider the common case of tagged objects with 3-bit tags and, for the sake of the example, assume that the tag 001 is chosen for floats. This setup was previously illustrated in Figure 1 where a tagged pointer refers to the heap-allocated float 3.02. The IEEE754 encoding of 3.02 happens to end with 001, the tag for floats in this example. This purely coincidental fact is bound to happen with $1/8$ of floating-point representations, since eight distinct tags exist on 3 bits.
Self-tagging exploits such occurrences where the tag bits of a pointer appear in its value. Such objects can be unboxed, making them tagged values instead of heap allocated value. However, since only $1/8$ of all floats can be unboxed in such a way, a second tag must be reserved for the remaining floats, which still need to be represented as heap allocated values with either tagged or generic pointers.
@@ 3. A First Experiment
This section presents the first experiment of float self-tagging with the Bigloo Scheme compiler (Manuel Serrano, 2024). Bigloo uses tagged objects by default and can be configured to use NaN-tagging (discussed in Section 4.2), making it ideal for comparing these techniques and self-tagging.
@@ 3.1. Overview of the Bigloo Compiler
The result of this compilation technique is that most temporary and local functions are precisely typed and avoid polymorphic representations, which contributes to removing otherwise necessary heap allocations and casts for floats. The compiler also tracks cases where only floats are stored into vectors, which are then transformed into vectors of C doubles. The experiment conducted in this section evaluates the impact of removing the memory allocations of floats that the compiler has not been able to remove statically. For a compiler that does not implement such static analyses, the impact might be even more significant than measured with Bigloo.
@@ 3.2. Experimental Setup
Both implementations are tested using a subset of the R7RS benchmarks, which is the standard benchmark suite for Scheme (r7r, 2024). The experiment includes all macro-benchmarks (more than 500 lines of code) and benchmarks whose calculations mainly involve floats (tagged as such in the R7RS benchmarks suite). For the purpose of analyzing results, it is useful to distinguish between float-intensive benchmarks (fibfp, fft, mbrot, nucleic, pnpoly, ray, simplex, sum1, and sumfp) and other benchmarks that use few of no floats.
The following results were obtained on a machine with an Intel Xeon W-2245 CPU 3.90GHz and 32 GB of RAM, under Debian 13 with Linux Kernel 6.10.6.

### v2

@@ Abstract.
Dynamic and polymorphic languages must attach information, such as types, to run time objects, and therefore adapt the memory layout of values to include space for this information. It makes it difficult to efficiently implement IEEE754 floating-point numbers as this format does not leave an easily accessible space to store the type information. The two main floating-point number encodings in-use to this day, tagged pointers and NaN-boxing, have drawbacks. Tagged pointers entail a heap-allocation of all float objects, and NaN-boxing puts additional run time costs on type checks and the handling of all other objects.
This paper presents self-tagging, a new approach to object tagging that can attach easily accessible type information to N-bit objects while retaining the ability to use all of their N bits for data. At its core, self-tagging exploits the fact that some bit sequences appear with high probability. Superimposing tags with these frequent sequences allows encoding both N-bit data and type within a single N-bit machine word. The main application of this approach is to represent IEEE754 64-bit and 32-bit floating point numbers on 64-bit and 32-bit machines respectively. We have implemented related variants of self-tagging in one JavaScript compiler and two distinct Scheme compilers to analyze their performance and compare them to tagged pointers and NaN-boxing. Our experiments demonstrate that in practice the approach eliminates heap-allocation of IEEE754 floating-point numbers and improves the execution time of float-intensive benchmarks in Scheme by 2.4$\times$, and in JavaScript by 3.6$\times$, with a negligible performance impact on other benchmarks, which makes it a good alternative to both tagged pointers and NaN-boxing.
Dynamic Languages, Polymorphic Languages, Floating-Point, NaN-Boxing, Compiler Optimization, JavaScript, Scheme
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Object-Oriented Programming, Systems, Languages & Applications; October 12–18, 2025; Singaporeisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Source code generationccs: Software and its engineering Polymorphism
@@ 1. Introduction
In dynamic and other polymorphic languages, efficiently handling floating-point numbers (floats) remains an elusive problem in search of an optimal solution. These languages require attaching type information to values at run time, including numeric types like integers and floats. This requirement conflicts with the IEEE754 standard (iee, 2019) for encoding floats, which uses a full 64-bit word to represent a double-precision float, leaving no space for additional type information. Various representations have been developed to mitigate this issue, each with its own trade offs.
This paper presents a novel technique for encoding floats that avoids most float allocations without adversely affecting the cost of checking types and accessing objects, which are frequent operations of dynamic languages. This technique does not require any program static analysis and could be integrated in most runtime systems of languages that need run time types. It is developed in the context of 64-bit runtime systems but also has applications on 32-bit systems. The rest of this section presents the main encodings in-use as well as their trade offs.
@@ 1.1. Tagged Objects
The classical and popular solution to preserve type information is to attach a tag to all objects (Gudeman, 1993; Leijen, 2022; Jones et al., 2012a; Appel, 1998). An N-bit machine word is used to encode an object reference (either a value or a pointer to a heap-allocated value) and a small bit field is reserved in the word to store type information. For the rest of the paper the term object is used interchangeably with object reference. Tagging allows low-cost type checks and object access at the cost of losing a few bits for data. Aligning all heap-allocated values to 64-bit machine words conveniently frees the low bits of pointers to store a 3-bit tag. On many architectures, accessing the fields of the object can be done at no additional cost by using an offset in the memory dereference instruction to take into account the tag for that type of object. Objects can thus be encoded using one of the three following representations.
@@
  • Generic pointers represent all remaining object types as pointers with a generic tag. Since the generic tag is associated with multiple types, type information is stored in the heap, typically in a header.
Hence, a tag indicates a type, but also a memory representation. A tagged value representation is the most efficient as it needs no heap allocation and no memory read. A tagged pointer still allows efficient type checks, but requires heap allocation. This introduces the overhead of dereferencing the pointer, and adds a strain on the garbage collector. A generic pointer is the least efficient representation as it involves storing type information in the heap, which increases space usage and the cost of type checks and memory management. Consequently, specific tags are a precious resource that must be carefully assigned to the most frequent types observed in programs, such as small integers or floats. Figure 1 shows a typical memory layout of each representation.
@@
As an example, consider an implementation using 3-bit tags. Due to the frequent use of small integers, a tagged value representation is appropriate for small integers and the choice of the tag 000 allows addition/subtraction directly on the tagged values. Since 3 bits are occupied by the tag, small integers are limited to 61 bits. This is generally considered an acceptable trade off since a two’s complement representation of signed integers can accommodate any number of bits, and 61 bits still offers a wide range of values. When larger integers are required, generic pointers to heap-allocated big integers can be used (Leijen, 2022). Consequently this provides a hybrid representation where the most common integers appear as efficient tagged values and less-frequent big integers are represented with more costly generic pointers.
Unfortunately, when it comes to the representation of floats, none of the aforementioned options are well-suited. Double-precision floats cannot be represented as tagged values since the IEEE754 standard enforces the use of 1 sign bit, 11 exponent bits and 52 mantissa bits (see Figure 2(a)), totaling 64 bits and leaving no space for a tag (iee, 2019). Parting from this standard is impracticable due to the set of instructions offered by x86 (Corporation, 2024) and ARM (Limited, 2009) architectures being specific to 32 or 64-bit floats.
@@
Still, tagged objects are straightforward to implement and are thus found in numerous compilers such as V8 (Sheludko and Solanes, 2020), QuickJS (Bellard and Gordon, 2024), and Hopc (Serrano, 2021) (JavaScript), the Lua interpreter (Roberto Ierusalimschy, 2024) (Lua), CRuby (Shaughnessy, 2013) (Ruby), SBCL (Newman et al., 2024) (Lisp), and Gambit (Marc Feeley, 2025) and Bigloo (Manuel Serrano, 2024) (Scheme). Recently, CPython also moved toward tagged pointers in the process of removing its global interpreter lock (Gross, 2023).
1.2. NaN-boxing
NaN-boxing circumvents the drawbacks of tagged pointer floats by reclaiming unused bits in the encodings of floating-point NaN values to store data (Gudeman, 1993; Leijen, 2022). As per the IEEE754 standard, a NaN value is represented by setting all 11 bits of the exponent to 1 and a non-zero mantissa (zero is used to encode Infinity). The mantissa’s highest bit distinguishes between a quiet and signaling NaN, which determines whether the NaN should signal an exception or fall through operations.
A subset of NaN’s can thus be reserved to encode non-float objects. A convenient subset is that of negative, quiet NaN’s, illustrated in Figure 2(b), which correspond to the interval from 0xfff8_0000_0000_0000 to 0xffff_ffff_ffff_ffff. This partitions floating-point numbers in two intervals. Bit patterns above 0xfff8_0000_0000_0000 are reserved for non-float objects. The bit patterns below 0xfff8_0000_0000_0000 are reserved for floats (including signaling NaN but not quiet NaN) and negative quiet NaN are mapped to the bit pattern 0xfff8_0000_0000_0000.
@@ 1.2. NaN-boxing
(b) NaN-boxing uses the 51 free bits on negative, quiet NaN values to store data.
@@
On current hardware, memory addresses typically fit in 48 bits. The 51 bits uncovered by NaN-boxing are thus sufficient for storing tagged pointers with 3-bit tags, and 32-bit small integers. It offers the advantage of unboxed floats, but impacts the performance of all other, non-float objects due to higher-cost machine instructions necessary to check the type and dereference NaN-boxed pointers.
This reliance on hardware specific details also interferes with other optimizations and portability. Moreover, NaN-boxing on 32-bit architectures would severely cripple memory management because 32-bit floats have NaNs that leave space for only 22-bit pointers, which only allows addressing 4 MiB.
NaN-boxing is used in a few language implementations, including SpiderMonkey (Foundation, 2024) (JavaScript), tinylisp (Engelen, 2024) (Lisp), LuaJIT (Pall, 2024) (Lua), and Zag (Mason, 2022) (Smalltalk).
1.3. Pointer-Biased NaN-boxing
NuN-boxing is a variant of NaN-boxing that alleviates the cost of dereferencing NaN-boxed pointers (Leijen, 2022). It relies on the (currently valid) fact that common hardware never returns negative, quiet NaN’s greater than or equal to 0xfffe_0000_0000_0000.
Figure 3. Encoding of Nun-boxed values, after having applied the 0x0001_0000_0000_0000 bias to floats. The low and high ranges (whose 16 highest bits are 0x0000 or 0xffff) are used for non-float objects that can be represented as tagged values or pointers.
Instead of taking advantage of the unused NaN value range from 0xfffe_0000_0000_0000 to 0xffff_ffff_ffff_ffff, which would result in an encoding similar to standard NaN-boxing, NuN-boxing biases all floats by adding 0x0001_0000_0000_0000. This frees up the lowest and highest ranges for 48-bit tagged objects as shown in Figure 3. It however reintroduces a cost for encoding and decoding floats (adding/subtracting the bias) to operate on their value.
NuN-boxing is used by JavaScriptCore (Inc., 2024) (JavaScript).
@@ 1.4. Contribution
The contribution of this paper is a thorough analysis of a new approach to object tagging, named self-tagging, that allows attaching type information to some 64-bit objects while retaining the ability to use all of their 64 bits for storing data. This approach is applied to implement double-precision floats as tagged values instead of tagged pointers. Contrary to NaN-boxing, the newly obtained representation does not impact the performance of encoding and decoding other tagged objects. It is also applicable in the context of single-precision floats on 32-bit architectures, which NaN-boxing does not support. Self-tagging cannot represent all floats but the conducted experiments show that the set of numbers it can represent is large enough to cover all floats actually used by all surveyed JavaScript and Scheme test programs. This paper shows that with float self-tagging there are very few if any heap-allocated floats in practice, thus providing an interesting trade off between the advantages of tagging and NaN-boxing.
@@ 1.5. Paper Structure
This paper is structured as follows. Section 2 describes the general idea of self-tagging with a focus on its application to floats. Section 3 discusses a basic form of self-tagging. Section 4 shows how to refine the approach to greatly reduce heap-allocation of floats. Section 5 presents technical details of its implementation. It shows portable implementations that could be used by compilers that generate either C or assembly code. Section 6 shows experimental results both in the context of Scheme with the Bigloo and Gambit compilers, and in the Hopc ahead-of-time JavaScript compiler. Section 7 discusses the adaptation of self-tagging to 32-bit architectures. Finally, section 8 presents related work.
@@ 2. Self-Tagging
This section describes the basic form of self-tagging, a tagging technique that exploits the fact that some values naturally contain the appropriate tag corresponding to their type, at the correct location, in their bit arrangement. While this may appear coincidental, it occurs frequently enough to allow the transformation of a significant proportion of heap allocated objects (either represented as tagged pointers or generic pointers) to tagged values. In this paper, unless stated otherwise, double-precision floats and a 64-bit architecture are implied. However, self-tagging pertains to 32-bit architectures as well, which is discussed in Section 7.
Consider the common case of tagged objects with 3-bit tags and, for the sake of the example, assume that the tag 001 is chosen for floats. This setup was previously illustrated in Figure 1 where a tagged pointer refers to the heap-allocated float 3.02. The IEEE754 encoding of 3.02 happens to end with 001, the tag for floats in this example. This purely coincidental fact is bound to happen with $1/8$ of all the bit patterns in the floating-point number space, since eight distinct tags exist on 3 bits.
Self-tagging exploits occurrences where the tag bits of a pointer appear in its value. Such objects can be unboxed, making them tagged values instead of heap allocated values. However, since only $1/8$ of all floats can be unboxed in such a way, a second tag must be reserved for the remaining floats, which still need to be heap allocated with either tagged or generic pointers.
@@ 3. A First Experiment
This section presents the first experiment of float self-tagging with the Bigloo Scheme compiler (Manuel Serrano, 2024). Bigloo uses tagged objects by default and can be configured to use NaN-boxing and NuN-boxing (discussed in Section 6.3), making it ideal for comparing these techniques and self-tagging.
@@ 3.1. Overview of the Bigloo Compiler
The result of this compilation technique is that most temporary and local functions are precisely typed and avoid polymorphic representations, which contributes to removing otherwise necessary heap allocations and casts for floats. The compiler also tracks cases where only floats are stored into vectors, which are then transformed into vectors of C doubles. The experiment conducted in this section evaluates the impact of removing the memory allocations of floats that the compiler is not able to remove statically. For a compiler that does not implement such static analyses, the impact might be even more significant than measured with Bigloo.
@@ 3.2. Experimental Setup
Both implementations are tested using a subset of the R7RS benchmarks, which is the standard benchmark suite for Scheme (r7r, 2024). The experiment includes all macro-benchmarks (more than 500 lines of code) and benchmarks whose calculations mainly involve floats (tagged as such in the R7RS benchmarks suite). For the purpose of analyzing results, it is useful to distinguish between benchmarks that use few or no floats and float-intensive benchmarks (fibfp=doubly recursive Fibonacci with floats, fft=1024 point Fast Fourier Transform, mbrot=mandelbrot set, nucleic=3D structure determination of a nucleic acid, pnpoly=determine if a 2D point is in a 20 sided polygon, ray=raytrace a scene with 33 spheres, simplex=linear programming using simplex method, sum1=add a file of floats, and sumfp=add the floats 0 to $10^{6}$).
The following results were obtained on a machine with an Intel Core i7-7700K, 48 GB of RAM, and under Debian 10.13 with kernel version SMP Debian 4.19.269-1.